OpenStudy (anonymous):

find the product checking to see if I am close (2r+4(2r-4) 4r^2-2r+8r

OpenStudy (phi):

is that (2r+4)(2r-4)? if so, the answer is not as posted. people use FOIL as a mnemonic, First , outer, inner last first terms 2r*2r outer 2r* -4 inner 4*2r last 4*-4 put it all together and simplify

OpenStudy (anonymous):

4r^2-6r is that close sorry im trying here

OpenStudy (anonymous):

because the 4*-4 cancel each other out correct

OpenStudy (phi):

first terms 2r*2r = 4r^2 outer 2r* -4 -8r inner 4*2r +8r last 4*-4 -16 (2r+4)(2r-4)= 4r^2-8r+8r-16 -8r+8r =0 (you have 8 r's and take away 8 r's) so the answer simplifies to 4r^2 - 16 notice that when you have the "same thing" except for + and - signs (a+b)(a-b) the middle terms cancel. you get a^2 - b^2 or in this case 4r^2 - 16

OpenStudy (anonymous):

ok so the final being the same thing cancels out the middle and the final answer is 4r^2-16 thank you for guiding me in the right direction this site rocks

OpenStudy (phi):

here is a short video that makes it clear http://www.khanacademy.org/math/algebra/polynomials/v/multiplying-binomials

OpenStudy (anonymous):

thank you